OpenStudy (anonymous):

Linear algebra: Express p(x) = 7 + 8x + 9x^2 as a linear combination of the following...

OpenStudy (anonymous):

\[p _{1}(x) = 2 + x 4x ^{2}\] \[p _{2}(x) = 1 - x + 3x ^{2}\] \[p _{3}(x) = 3 + 2x + 5x ^{?}\]